[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jul 21, 2016
nhc.tsbadmin at noaa.gov
Thu Jul 21 12:44:59 UTC 2016
Here is today's summary of NHC computer operations:
--- Craig Mattocks
1. Update on the overhead Lights in the TAFB ops area. William Martinez
will follow up. The light bulbs are good, but the light fixture over the
backup desk is bad and needs to be replaced.
2. Update on the security camera display. NHC Administrator Officer Jann
Perez reports: "We are in the process of upgrading our security system. As
a result, some of the security monitors will be down until the proper
connections are made by the contractor. Most of the new cameras are up and
running. Our downtime should be minimal."
1. ATCF - TSB dataflow expert Monica Bozeman has adjusted the ingest time
and frequency in the unix crontab entry for the TAFB fixes. The TAFB Dvorak
fixes are now importing on time.
2. ATCF - The FSU Superensemble arrived several hours late yesterday for
both Darby and Estelle at 12Z. Senior Hurricane Specialist Mike Brennan
will contact WeatherPredict Consulting, Inc. to ask what is causing this
3. N-AWIPS/AWIPS2 - NASA SPoRT microwave imagery stopped coming into
N-AWIPS at 10Z yesterday, but the problem seemed to have resolved itself by
4. ATCF - The Pacific ATCF workstation continues to be sluggish at times.
NCO SysAdmin Brian Maher is continuing to monitor the situation.
5. N-AWIPS/AWIPS-2 - The CPU load bar in the lower left of the display on
the Atlantic N-AWIPS workstation Ridges has been very high even when the
forecaster was doing very little work on the machine. Logging out and back
in had no effect, but there has been a gradual decrease in activity since
then. If this continues, TSB will submit a Help Desk ticket to NCO.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts